OpenStudy (iwanttogotostanford):

HELP W? AN ACT PREP QUESTION< PLEASE I NEED TO DO WELL: A ramp for wheelchair access to the gym has a slope of 5% (that is, the ramp rises 5 feet vertically for every 100 feet of horizontal distance). The entire ramp is built on level ground, and the entrance to the gym is 2 feet above the ground. What is the horizontal distance, in get, between the ends of the ramp???

OpenStudy (iwanttogotostanford):

@Vocaloid

OpenStudy (iwanttogotostanford):

@jim_thompson5910 @amistre64 @pooja195 @satellite73

jimthompson5910 (jim_thompson5910):

do you have a pic of the problem drawn out? or are you stuck on that part?

OpenStudy (iwanttogotostanford):

thats all of the problem and thats all it gives me

OpenStudy (iwanttogotostanford):

i have the answer choices but i really just need to know how to do it

jimthompson5910 (jim_thompson5910):

well I meant in terms of you drawing it yourself (not having the teacher draw it for you)

OpenStudy (iwanttogotostanford):

i have an ACT prep book but sure

jimthompson5910 (jim_thompson5910):

here's what the drawing would look like 

OpenStudy (iwanttogotostanford):

i just need to know the process, i already have the answer

OpenStudy (iwanttogotostanford):

oh ok thanks thats helpful

jimthompson5910 (jim_thompson5910):

the info `A ramp for wheelchair access to the gym has a slope of 5% (that is, the ramp rises 5 feet vertically for every 100 feet of horizontal distance)` basically tells us that we'll have this proportion 2/x = 5/100

jimthompson5910 (jim_thompson5910):

the numerators represent the vertical change in distance the denomiators represent the horizontal change in distance

OpenStudy (iwanttogotostanford):

ok

jimthompson5910 (jim_thompson5910):

so you see how I got 2/x = 5/100? if so, then the next step is to isolate x

OpenStudy (iwanttogotostanford):

yep, I got that

jimthompson5910 (jim_thompson5910):

ok tell me what x is

OpenStudy (iwanttogotostanford):

40

jimthompson5910 (jim_thompson5910):

good

OpenStudy (iwanttogotostanford):

so thats the answer and how we do it?

jimthompson5910 (jim_thompson5910):

like I said, solving 2/x = 5/100 for x

OpenStudy (iwanttogotostanford):

ok, thanks! I have another question! can i tag you in a new question and help me?

jimthompson5910 (jim_thompson5910):

2/x = 5/100 2*100 = x*5 ... cross multiply 200 = 5x 5x = 200 5x/5 = 200/5 ... dividing both sides by 5 x = 40

OpenStudy (iwanttogotostanford):

ok! Can i make a new question and tag you in it?

jimthompson5910 (jim_thompson5910):

sure